OpenStudy (anonymous):

|f| – 2/3 = 5/6 ------------- 5/6 or -5/6 2/3 or -2/3 1 1/2 or -1 1/2 (one and a half) no solution

OpenStudy (anonymous):

THANK YOU FOR ANSWERING!!!!!!!! TOUGH STUFF!

OpenStudy (aakashsudhakar):

So you're looking to solve for f. The first thing you need to do is to get everything into equal terms. |f| - (2/3) = (5/6) |f| - (4/6) = (5/6) f may be in absolute value notation, but (4/6) sure isn't. You can definitely isolate f easily. |f| = 9/6 or... |f| = 1 + (3/6) or even... |f| = 1 + (1/2) Since it's absolute value notation, this translates to... f = ±[1 + (1/2)] This matches up with the third option, which I'm just going to call (C).

OpenStudy (anonymous):

HI